IMN2) Breakdowns and defective products produced for the last 15 days on a CNC machine are presented in the following table: . 1 2 3 4 5 6 7 8 9 10 11 12 13 14 15 Day Breakdown Breakdown Breakdown code AB B C D DEF G start time 3:00-16:00-10:00 17:00 6:00 15:00 3:00 end time 5:00一 一 2:00 11:00-3:00 12:00- 21:00 8:00 . The breakdown code represents each recorded breakdown. If the same code appears on more tharn one day, it is because the repair did not take place or was not finished on the preceding day. The factory works a. Calculate the machine statistical availability beginning at 0:00 hour on day 1 and finishing at 0:00 hour on day 16 (or 24:00 on day 15). b. What is the difference between statistical availability and the overall equipment effleieney (OEE) rate? . Briefly explain
Answer #1

a. Total breakdown = sum of all breakdowm times ( A to G)= 40 hours.

Total working hours = 24x15 = 360

Availability = 360-40 = 320 hours

Statistical availability = run time / total production time = 320 /360 = 88.88%

b. OEE is a broader measure where three counts, viz. availabilty, performance and quality are taken into account.

Performance = net run time / run time = ideal run time x total count / run time

Quality = Good count / total count

OEE = Availability x Performance x Quality

To determine OEE, in addition to the down time Vs. total scheduled time, inputs on quality and performance are also needed.